OpenStudy (anonymous):

evaluate and simplify [f(x+h)-f(x)]/h if f(x)=x^2 -2x

OpenStudy (anonymous):

$$\frac{f(x+h)-f(x)}h=\frac{(x+h)^2-2(x+h)-x^2+2x}h\\\ \ \ =\frac{x^2+2hx+h^2-2x-2h-x^2+2x}h\\\ \ \ =\frac{h^2+2hx-2h}h=2x-2+h$$

OpenStudy (anonymous):

@Skaternerd123 smells take a shower kid

OpenStudy (anonymous):

@Skaternerd123 is in 4th grade for real

OpenStudy (anonymous):

thanks!!!!